Z is a standard normal variable. Compute the requested probability when
P(-.87 ≤Z≤ 2.21)

Found 2 solutions by Boreal, jim_thompson5910:
Answer by Boreal(15235)

This can be done from a table or a calculator.
From the calculator use 2nd VARS 2 and (-0.87,2.21)
probability is 0.7943.

Answer by jim_thompson5910(35256)


I'm going to use this table (a similar one should be found in the back of your textbook)

Using that table, we can find the following:
P(Z < -0.87) = 0.1922 (see this image to see how I got the value)
P(Z < 2.21) = 0.9864 (see this image to see how I got the value)

So,
P(a < Z < b) = P(Z < b) - P(Z < a)
P(-0.87 < Z < 2.21) = P(Z < 2.21) - P(Z < -0.87)
P(-0.87 < Z < 2.21) = 0.9864 - 0.1922
P(-0.87 < Z < 2.21) = 0.7942 (which is the final answer)
The values are approximate